The 84,000 million liters of water saved in Sydney, Australia, in 1 year represents 400 more million liters than double the amount saved in 2006. How much water was saved in Sydney, Australia, in 2006? (I think I might have to use th formula: W(t)=2tsquared-13t+39)
Answer by Theo(13342)

equation is:
84,000 = 2*x + 400
we're dealing in millions of liters.
subtract 400 from both sides of this equation to get:
83,600 = 2*x
divide both sides of this equation to get:
41,800 = x
41,800 million liters of water were saved in 2006.
this year, you saved 400 million liters more than 2 times that amount.
that means that this year you saved 2 * (41,800) + 400 = 83,600 + 400 = 84,000 million liters of water.